# The Cohen–Lenstra moments over function fields via the stable homology of non-splitting Hurwitz spaces

### v2

@@ 5.3. Applying the lemma to count points on Hurwitz spaces
This follows from § 5.2, taking the constant $C^{\prime}$ there to be the same as $C$, once we verify its hypotheses. We will take $\ell>\max(n,|G|,q)$. To simplify matters slightly, we will assume all irreducible components of $Y_{n}$ are geometrically irreducible, as components which are not geometrically irreducible have no $\mathbb{F}_{q}$ points, by smoothness of the $Y_{n}$. Let $B$ denote the spectrum of a henselian dvr with residue field $\mathbb{F}_{q}$ and generic point of characteristic 00. In this case, there is a scheme $Z_{n}$ which is a union of $r$ irreducible components of $\operatorname{CHur}^{G,c,g}_{n,B}$, so that $Y_{n}=Z_{n}\times_{B}\operatorname{Spec}\mathbb{F}_{q}$. We let $Y_{n,\mathbb{C}}:=Z_{n}\times_{B}\operatorname{Spec}\mathbb{C}$ and $Y_{n,\overline{\mathbb{F}}_{q}}:=Y_{n}\times_{\operatorname{Spec}\mathbb{F}_{q}}\operatorname{Spec}\overline{\mathbb{F}}_{q}.$
We first verify the stability assumption from § 5.2(1). To do this, we start by constructing the vector spaces $V_{i}$, with the Frobenius action. For $B$ a scheme, we use $\operatorname{Conf}_{n,B}$ denote the configuration space parameterizing $n$ unordered points on $\mathbb{A}^{1}_{B}$. (Previously, we had used $\operatorname{Conf}_{n}$ as notation for this when $B=\operatorname{Spec}\mathbb{C}$.) Now, define $V_{i}:=H^{i}(\operatorname{Conf}_{n,\overline{\mathbb{F}}_{q}};\mathbb{Q}_{\ell})^{\oplus r}$ as a vector space with Frobenius action, where $r$ is the number of geometrically irreducible components of $Y_{n}$ for sufficiently large $n$. Note that the dimension of $V_{i}$ is independent of $n$ and $q$ once $n>2$ as a vector space with Frobenius action, and the Frobenius action is given by and identity matrix of dimension $r$ if $i=0$, $q$ times the identity matrix of dimension $r$ when $i=1$, and the 00 matrix if $i>1$. In particular, we have $\sum_{i=0}^{\infty}(-1)^{i}\operatorname{tr}(\operatorname{Frob}_{q}^{-1}|V_{i})=r(1-1/q)$.
To verify § 5.2(1), we next need to show the trace of Frobenius on $V_{i}$ agrees with the trace of Frobenius on $H^{i}(Y_{n,\overline{\mathbb{F}}_{q}};\mathbb{Q}_{\ell})$ for $n>iI+J$. There is a map
@@
which is an isomorphism for $n>iI+J$ by Theorem 1.4.5. It follows from [EVW16, Proposition 7.7] that there is an isomorphism
@@
once we exhibit a suitable normal crossings compactification of configuration space. This is afforded by [EL23, Corollary B.1.4]. Therefore, (5.6) restricts to an isomorphism
@@
Combining (5.5), (5.7), and (5.8), we obtain an Frobenius equivariant isomorphism $U_{i}:H^{i}(\operatorname{Conf}_{n,{\overline{\mathbb{F}}_{q}}};\mathbb{Q}_{\ell})^{\oplus r}\to H^{i}(Y_{n,\overline{\mathbb{F}}_{q}};\mathbb{Q}_{\ell})$. This implies $\operatorname{tr}(\operatorname{Frob}_{q}|V_{i})=\operatorname{tr}(\operatorname{Frob}_{q}|H^{i}(Y_{n,\overline{\mathbb{F}}_{q}};\mathbb{Q}_{\ell}))$.
Finally, we verify the exponential bound assumption, § 5.2(2). This condition with $C=C^{\prime}$ nearly follows from [EVW16, Proposition 7.8], using our assumption that $\ell>\max(n,|G|,q)$ as well as the comparison between étale and Betti cohomology. The slight difference we need is that [EVW16, Proposition 7.8] is proven for Hurwitz stacks, instead of pointed Hurwitz spaces. However, the same proof works for pointed Hurwitz spaces, since the analog of [EVW16, Corollary 6.2] without quotienting by the $G$ conjugation action can be deduced in the same way from [EVW16, Theorem 6.1]. We note that we can take the value of $C$ to be independent of $q$ and $\ell$ satisfying our constraints because $\dim H^{i}(Y_{n},\mathbb{Q}_{\ell})\leq\dim H^{i}(\operatorname{CHur}^{G,c,g}_{n,\mathbb{C}};\mathbb{Q}_{\ell})$ is independent of $q$ and $\ell$. ∎
@@
Once we verify the claim about the number of geometrically irreducible components, (5.9) follows from § 5.3. The remainder of the proof is essentially explained in [SW23, Theorem 3.1], and we now spell out the details. In the case that $n$ is even, [LWZB24, Lemma 10.3] yields a bijection between components of $\operatorname{CHur}^{G,c,g}_{n,\overline{\mathbb{F}}_{q}}$ and components of $\operatorname{CHur}^{G,c,g}_{n,\mathbb{C}}$. Then, [LWZB24, Theorem 12.1(2)] and [LWZB24, Corollary 12.6] yield that the number of irreducible components of $\operatorname{CHur}^{G,c,g}_{n,\mathbb{C}}$ can be identified bijectively with a torsor for the set of elements of a certain finite group $H_{2}(G,c)$ which are equal to their $q$th power. There is an identification $H_{2}(G,c)\simeq(\wedge^{2}H)^{\mathbb{Z}/2\mathbb{Z}}$ as is explained in the course of the proof of [SW23, Theorem 3.1]. Moreover, in our case, $\mathbb{Z}/2\mathbb{Z}$ acts on $H$ by inversion, and hence it acts trivially on $\wedge^{2}H$, and so $(\wedge^{2}H)^{\mathbb{Z}/2\mathbb{Z}}\simeq\wedge^{2}H$. In other words, the number of geometrically irreducible components is precisely the $q-1$ torsion subgroup of $\wedge^{2}H$, $\wedge^{2}H[q-1]$. Since $h:=\gcd(|H|,q-1)$, this is also equal to $\wedge^{2}H[h]$, and hence there are $\left|\wedge^{2}H[h]\right|$ geometrically irreducible components of $\operatorname{CHur}^{G,c,g}_{n,{\mathbb{F}}_{q}}$.
@@ 5.5. Non-abelian Cohen–Lenstra moments
where $[H:H^{\mathbb{Z}/2\mathbb{Z}}]$ denotes the index of the $\mathbb{Z}/2\mathbb{Z}$ invariants of $H$ in $H$.
@@
We claim there is a unique geometrically irreducible component of $\operatorname{CHur}^{G,c,\operatorname{id}}_{n,\mathbb{F}_{q}}$, whose $\omega$-invariant is $\delta$ precisely when $\operatorname{ord}(\operatorname{im}\delta)\mid q-1$ and $n$ is even, where the $\omega$-invariant of a component was defined (and shown to be well defined) in [Liu22, Proposition 4.3]. To prove this claim, we note that $\operatorname{CHur}^{\mathbb{Z}/2\mathbb{Z},\iota}_{n,\mathbb{F}_{q}}$ has a unique geometrically irreducible component, where $\iota$ is the nontrivial element of $\mathbb{Z}/2\mathbb{Z}$. The claim then follows from [Liu22, Lemma 4.4], [Liu22, Lemma 4.5], and [LWZB24, Proposition 12.7]. Here, we are also using that there is a unique component of the Hurwitz scheme parameterizing quadratic extensions of $\mathbb{F}_{q}(t)$ of reduced discriminant $q^{n}$ precisely when $n$ is even, as follows from Riemann Hurwitz, and so the $M_{a}$ appearing in [Liu22, Lemma 4.4] can all be taken to be $2$ and the $E_{a}$ can be taken to be the singleton 00, using the equality in [Liu22, Lemma 4.5].
The result now follows from (5.4) with $r=1$ and [Liu22, Lemma 4.6], which accounts for the factor of $[H:H^{\mathbb{Z}/2\mathbb{Z}}]$. ∎
@@ A.2. A monoid model of Hurwitz spaces
Define the topological space $\mathrm{conf}$ as the space whose points are given by pairs $(t,x)$ where, $t\in\mathbb{R}_{\geq 0}$, and $x$ is a (possibly empty) configuration of finitely many distinct unordered points $(0,t)\times(0,1)$, not intersecting $[0,\frac{1}{2})\times[0,1]$ or $(t-\frac{1}{2},t]\times[0,1]$. The multiplication is given by concatenation of configurations and adding the elements $t$. We use $\mathrm{conf}$ to denote this topological monoid, and $\mathrm{conf}_{n}$ to denote the component with configuration of $n$ points. There is a map of topological monoids $t:\mathrm{conf}\to\mathbb{R}_{\geq 0}$ given by $(t,x)\mapsto t$. There is also a subset $\mathrm{ord}\subset\mathrm{conf}$ consisting of configurations such that the second coordinate of each point in $x$ is $\frac{1}{2}$. For each $n$, the intersection $\mathrm{ord}\cap\mathrm{conf}_{n}$ is contractible. We will use the inclusion $\mathrm{ord}_{n}:=\mathrm{ord}\cap\mathrm{conf}_{n}\subset\mathrm{conf}_{n}$ to view each component of $\mathrm{conf}$ as being equipped with a ‘base point’, i.e a fixed contractible subspace. We also use $\mathrm{confbig}$ to denote the variant of this topological monoid where we remove the condition that configurations not intersect $[0,\frac{1}{2})\times[0,1]$ and $(t-\frac{1}{2},t]\times[0,1]$, and use $\mathrm{bigord}$ and $\mathrm{bigord}_{n}$ to denote the corresponding subspaces. There is an inclusion of topological monoids $\mathrm{conf}\to\mathrm{confbig}$ which is clearly a homotopy equivalence.
@@
By convention, the composition in the fundamental group/groupoid of a space is such that given two paths $\gamma$, $\gamma^{\prime}$, we have $[\gamma][\gamma^{\prime}]$ is the class of the path which is first $\gamma^{\prime}$ and then $\gamma$ (assuming they are composable). This convention is the one that is compatible with the orientation of the interval in that we view $\gamma^{\prime}$ as a morphism from its starting point to its end point.
@@
Since $\operatorname{Hur}^{X}$, viewed as an $\mathbb{E}_{1}$-algebra, is obtained as a multiplicative finite cover of $\operatorname{Conf}$, we can construct a model $\operatorname{hur}^{X}$ for $\operatorname{Hur}^{X}$ as the corresponding multiplicative finite cover of $\mathrm{conf}$ in topological spaces. Namely, we first use the base points from the monoid map $\mathbb{N}\to\mathrm{conf}$ constructed in § A.2 to identify the fundamental group of $\mathrm{conf}$ with $B_{n}$, using the convention of § A.2. Then, $\operatorname{hur}^{X}_{n}$, as a cover of $\mathrm{conf}_{n}$, can be constructed as the quotient of $X^{n}\times\widetilde{\mathrm{conf}}_{n}$ by the braid group $B_{n}$, where $\widetilde{\mathrm{conf}}_{n}$ is the universal cover of $\mathrm{conf}_{n}$; More explicitly, a point of $\operatorname{hur}^{X}_{n}$ is given by a $B_{n}$-equivalence class of data $(x,t,\gamma,\alpha_{i})$, where $(x,t)\in\mathrm{conf}_{n}$, $\gamma$ is a homotopy class of paths from this point to $\mathrm{ord}$, and $\alpha_{i}$, for $1\leq i\leq n$, is an $n$ tuple of points in $X$, where $n=|x|$. Here the braid group action is such that $j$th standard generator of $B_{n}$, $\sigma_{j}$, sends $(x,t,\gamma,\alpha_{i})$ to $(x,t,\sigma_{j}\gamma,\sigma_{j}(\alpha_{i}))$ using § A.2 and § A.2 to get a left action of $B_{n}$ on such $\gamma$, and using the left action of the braid group on $X^{n}$ from § A.1.
@@
The key construction we use for Hurwitz spaces for braided sets is the following cutting construction. The idea for this construction is that if we start with a configuration in $[0,t^{\prime}]\times[0,1]$ which never contains points in the vertical line $t\times[0,1]$, we can cut it into two configurations by cutting along $t\times[0,1]$.
@@
Let $t\in[0,\infty)$, and consider the open set $U(t)$ of $\operatorname{hurbig}^{X}$ consisting of $(x,t^{\prime},\gamma,\alpha_{i})$ such that $t^{\prime}>t$ and no points of the configuration $x$ lie on $t\times[0,1]$. Then there is a continuous map $U(t)\to\operatorname{hurbig}^{X}\times\operatorname{hurbig}^{X}$ which we call ‘cutting at $t$’, given as follows: We first cut the configuration on $[0,t^{\prime}]$ into two configurations $x^{\prime},x^{\prime\prime}$ on $[0,t]$ and $[0,t^{\prime}-t]\cong[t,t^{\prime}]$. Choose a homotopy class of paths $\phi$ from $x$ to $\mathrm{ord}$ with a representative that doesn’t cross $t\times[0,1]$. Using the braid group action, the pair $(x,t^{\prime},\gamma,\alpha_{i})$ is equivalent to $(x,t^{\prime},\phi,\beta_{i})$ for some $\beta_{i}$ obtained by acting on the tuple $(x_{i})$ via the element of the braid group given by the class of $\phi\gamma^{-1}$. We can now similarly cut $\phi$ into homotopy classes of paths $\phi^{\prime},\phi^{\prime\prime}$ from the configurations $x^{\prime},x^{\prime\prime}$ to the respective base points, and cut the sequence $\beta_{i}$ into $\alpha_{j}^{\prime},\alpha_{k}^{\prime\prime}$ by unconcatenating it into two sequences corresponding to the number of points in the configuration in $[0,t]\times[0,1]$ and $[t,t^{\prime}]\times[0,1]$ respectively.
Then the cutting map sends $(x,t^{\prime},\gamma,\alpha_{i})$ to $(x^{\prime},t,\phi^{\prime},\alpha^{\prime}_{j})$ and $(x^{\prime\prime},t^{\prime}-t,\phi^{\prime\prime},\alpha^{\prime\prime}_{k})$. Since the concatenation map $X^{a}\times X^{b}\to X^{a+b}$ is $B_{a}\times B_{b}$-equivariant, this doesn’t depend on the choice of $\phi$.

### v3

@@ 5.3. Applying the lemma to count points on Hurwitz spaces
This follows from § 5.2, taking the constant $C^{\prime}$ there to be the same as $C$, once we verify its hypotheses.
To simplify matters slightly, we will assume all irreducible components of $Y_{n}$ are geometrically irreducible, as components which are not geometrically irreducible have no $\mathbb{F}_{q}$ points, by smoothness of the $Y_{n}$. Let $B$ denote the spectrum of a henselian dvr with residue field $\mathbb{F}_{q}$ and generic point of characteristic 0. In this case, there is a scheme $Z_{n}$ which is a union of $r$ irreducible components of $\operatorname{CHur}^{G,c,g}_{n,B}$, so that $Y_{n}=Z_{n}\times_{B}\operatorname{Spec}\mathbb{F}_{q}$. We let $Y_{n,\mathbb{C}}:=Z_{n}\times_{B}\operatorname{Spec}\mathbb{C}$ and $Y_{n,\overline{\mathbb{F}}_{q}}:=Y_{n}\times_{\operatorname{Spec}\mathbb{F}_{q}}\operatorname{Spec}\overline{\mathbb{F}}_{q}.$
We first verify the stability assumption from § 5.2(1). To do this, we start by constructing numbers $t_{i}$, which we will realize as the trace of $\operatorname{Frob}_{q}^{-1}$ acting on certain vector spaces. For $B$ a scheme, we use $\operatorname{Conf}_{n,B}$ denote the configuration space parameterizing $n$ unordered points on $\mathbb{A}^{1}_{B}$. (Previously, we had used $\operatorname{Conf}_{n}$ as notation for this when $B=\operatorname{Spec}\mathbb{C}$.) Choose $\ell$ prime to $|G|$ and $q$. Define $t_{i}:=\operatorname{tr}(\operatorname{Frob}_{q}^{-1}|H^{i}(\operatorname{Conf}_{n,\overline{\mathbb{F}}_{q}};\mathbb{Q}_{\ell}))^{\oplus r}$ as a vector space with Frobenius action, where $r$ is the number of geometrically irreducible components of $Y_{n}$ for sufficiently large $n$. Note that $t_{i}$ is independent of $n$ and $q$ once $n>2$ and is given by the trace of the identity matrix of dimension $r$ if $i=0$, the trace of $q$ times the identity matrix of dimension $r$ when $i=1$, and 0 if $i>1$. In particular, we have $\sum_{i=0}^{\infty}(-1)^{i}t_{i}=r(1-1/q)$.
To verify § 5.2(1), we next need to show the $t_{i}$ agrees with the trace of Frobenius on $H^{i}(Y_{n,\overline{\mathbb{F}}_{q}};\mathbb{Q}_{\ell})$ for $n>iI+J$. There is a map
@@
which is an isomorphism for $n>iI+J$ by Theorem 1.4.5.
Note that configuration space has a normal crossings compactification by [EL23, Corollary B.1.4]. It follows from [EVW16, Proposition 7.7] that for $\ell$ prime to $|G|$ and $q$ and $j\geq 1$ that there is an isomorphism
(5.6) $H^{i}(Y_{n,\overline{\mathbb{F}}_{q}},\mathbb{Z}/\ell^{j}\mathbb{)}\simeq H^{i}(Y_{n,\mathbb{C}},\mathbb{Z}/\ell^{j}\mathbb{Z}).$
Technically speaking, [EVW16, Proposition 7.7] assumed $j=1$, but the same proof works for $j>1$. Hence taking the limit over $j$ and tensoring with $\mathbb{Q}_{\ell}$, we obtain
@@
when $\ell$ is prime to $|G|$ and $q$.
Therefore, (5.7) restricts to an isomorphism
@@
Combining (5.5), (5.8), and (5.9), we obtain an isomorphism $U_{i}:H^{i}(\operatorname{Conf}_{n,{\overline{\mathbb{F}}_{q}}};\mathbb{Q}_{\ell})^{\oplus r}\to H^{i}(Y_{n,\overline{\mathbb{F}}_{q}};\mathbb{Q}_{\ell})$. This isomorphism $U_{i}$ is moreover Frobenius equivariant because it is induced by a map of schemes $Y_{n}\to\coprod_{i=1}^{r}\operatorname{Conf}_{n,\mathbb{F}_{q}}$ over $\mathbb{F}_{q}$. This implies $t_{i}=\operatorname{tr}(\operatorname{Frob}_{q}|H^{i}(Y_{n,\overline{\mathbb{F}}_{q}};\mathbb{Q}_{\ell}))$.
Finally, we verify the exponential bound assumption, § 5.2(2). This condition with $C=C^{\prime}$ follows from an argument similar to [EVW16, Proposition 7.8]. (It would exactly follow from [EVW16, Proposition 7.8], except the statement of that result assumes $\ell>n$, which we don’t want to assume. The reason they have that restriction is that they argue by passing to a cover where they order the $n$ branch points and so they need to invert $n!$. However, we can avoid passing to such a cover and directly compare the cohomology between in characteristic 0 and positive characteristic, so we do not need to invert $n!$.) Recall (5.7) holds, as we showed above. After identifying étale cohomology of $Y_{n,\mathbb{C}}$ with singular cohomology, the result then follows from [EVW16, Proposition 2.5].
We note that we can take the value of $C$ to be independent of $q$ and $\ell$ prime to $|G|$ and $q$ satisfying our constraints because $\dim H^{i}(Y_{n},\mathbb{Q}_{\ell})\leq\dim H^{i}(\operatorname{CHur}^{G,c,g}_{n,\mathbb{C}};\mathbb{Q}_{\ell})$ is independent of $q$ and $\ell$. ∎
@@
Once we verify the claim about the number of geometrically irreducible components, (5.10) follows from § 5.3. The remainder of the proof is essentially explained in [SW23, Theorem 3.1], and we now spell out the details. In the case that $n$ is even, [LWZB24, Lemma 10.3] yields a bijection between components of $\operatorname{CHur}^{G,c,g}_{n,\overline{\mathbb{F}}_{q}}$ and components of $\operatorname{CHur}^{G,c,g}_{n,\mathbb{C}}$. Then, [LWZB24, Theorem 12.1(2)] and [LWZB24, Corollary 12.6] yield that the number of irreducible components of $\operatorname{CHur}^{G,c,g}_{n,\mathbb{C}}$ can be identified bijectively with a torsor for the set of elements of a certain finite group $H_{2}(G,c)$ which are equal to their $q$th power. There is an identification $H_{2}(G,c)\simeq(\wedge^{2}H)^{\mathbb{Z}/2\mathbb{Z}}$ as is explained in the course of the proof of [SW23, Theorem 3.1]. Moreover, in our case, $\mathbb{Z}/2\mathbb{Z}$ acts on $H$ by inversion, and hence it acts trivially on $\wedge^{2}H$, and so $(\wedge^{2}H)^{\mathbb{Z}/2\mathbb{Z}}\simeq\wedge^{2}H$. In other words, the number of geometrically irreducible components is precisely the q1q-1 torsion subgroup of $\wedge^{2}H$, $\wedge^{2}H[q-1]$. Since $h:=\gcd(|H|,q-1)$, this is also equal to $\wedge^{2}H[h]$, and hence there are $\left|\wedge^{2}H[h]\right|$ geometrically irreducible components of $\operatorname{CHur}^{G,c,g}_{n,{\mathbb{F}}_{q}}$.
@@ 5.5. Non-abelian Cohen–Lenstra moments
where $H:H^{\mathbb{Z}/2\mathbb{Z}}]$ denotes the index of the $\mathbb{Z}/2\mathbb{Z}$ invariants of $H$ in $H$.
@@
We claim there is a unique geometrically irreducible component of $\operatorname{CHur}^{G,c,\operatorname{id}}_{n,\mathbb{F}_{q}}$, whose $\omega$-invariant is $\delta$ precisely when $\operatorname{ord}(\operatorname{im}\delta)\mid q-1$ and $n$ is even, where the $\omega$-invariant of a component was defined (and shown to be well defined) in [Liu22, Proposition 4.3]. To prove this claim, we note that $\operatorname{CHur}^{\mathbb{Z}/2\mathbb{Z},\iota}_{n,\mathbb{F}_{q}}$ has a unique geometrically irreducible component, where $\iota$ is the nontrivial element of $\mathbb{Z}/2\mathbb{Z}$. The claim then follows from [Liu22, Lemma 4.4], [Liu22, Lemma 4.5], and [LWZB24, Proposition 12.7]. Here, we are also using that there is a unique component of the Hurwitz scheme parameterizing quadratic extensions of $\mathbb{F}_{q}(t)$ of reduced discriminant $q^{n}$ precisely when $n$ is even, as follows from Riemann Hurwitz, and so the $M_{a}$ appearing in [Liu22, Lemma 4.4] can all be taken to be $2$ and the $E_{a}$ can be taken to be the singleton 0, using the equality in [Liu22, Lemma 4.5].
The result now follows from (5.4) with $r=1$ and [Liu22, Lemma 4.6], which accounts for the factor of $H:H^{\mathbb{Z}/2\mathbb{Z}}]$. ∎
@@ A.2. A monoid model of Hurwitz spaces
Define the topological space $\mathrm{conf}$ as the space whose points are given by pairs $(t,x)$ where, $t\in\mathbb{R}_{\geq 0}$, and $x$ is a (possibly empty) configuration of finitely many distinct unordered points in $(0,t)\times(0,1)$, not intersecting $0,\frac{1}{2})\times[0,1]$ or $(t-\frac{1}{2},t]\times[0,1]$. The multiplication is given by concatenation of configurations and adding the elements $t$. We use $\mathrm{conf}$ to denote this topological monoid, and $\mathrm{conf}_{n}$ to denote the component with configuration of $n$ points. There is a map of topological monoids $t:\mathrm{conf}\to\mathbb{R}_{\geq 0}$ given by $(t,x)\mapsto t$. There is also a subset $\mathrm{ord}\subset\mathrm{conf}$ consisting of configurations such that the second coordinate of each point in $x$ is $\frac{1}{2}$. For each $n$, the intersection $\mathrm{ord}\cap\mathrm{conf}_{n}$ is contractible. We will use the inclusion $\mathrm{ord}_{n}:=\mathrm{ord}\cap\mathrm{conf}_{n}\subset\mathrm{conf}_{n}$ to view each component of $\mathrm{conf}$ as being equipped with a ‘base point’, i.e a fixed contractible subspace. We also use $\mathrm{confbig}$ to denote the variant of this topological monoid where we remove the condition that configurations not intersect $0,\frac{1}{2})\times[0,1]$ and $(t-\frac{1}{2},t]\times[0,1]$, and use $\mathrm{bigord}$ and $\mathrm{bigord}_{n}$ to denote the corresponding subspaces. There is an inclusion of topological monoids $\mathrm{conf}\to\mathrm{confbig}$ which is clearly a homotopy equivalence.
@@
By convention, the composition in the fundamental group/groupoid of a space is such that given two paths $\gamma$, $\gamma^{\prime}$, we have $\gamma][\gamma^{\prime}]$ is the class of the path which is first $\gamma^{\prime}$ and then $\gamma$ (assuming they are composable). This convention is the one that is compatible with the orientation of the interval in that we view $\gamma^{\prime}$ as a morphism from its starting point to its end point.
@@
Since $\operatorname{Hur}^{X}$, viewed as an $\mathbb{E}_{1}$-algebra, is obtained as a multiplicative finite cover of $\operatorname{Conf}$, we can construct a model $\operatorname{hur}^{X}$ for $\operatorname{Hur}^{X}$ as the corresponding multiplicative finite cover of $\mathrm{conf}$ in topological spaces. Namely, we first use the base points constructed in § A.2 to define the monoid map $\mathbb{N}\to\mathrm{conf}$, which allows us to identify the fundamental group of $\mathrm{conf}$ with $B_{n}$, using the convention of § A.2. Then, $\operatorname{hur}^{X}_{n}$, as a cover of $\mathrm{conf}_{n}$, can be constructed as the quotient of $X^{n}\times\widetilde{\mathrm{conf}}_{n}$ by the braid group $B_{n}$, where $\widetilde{\mathrm{conf}}_{n}$ is the universal cover of $\mathrm{conf}_{n}$; More explicitly, a point of $\operatorname{hur}^{X}_{n}$ is given by a $B_{n}$-equivalence class of data $(x,t,\gamma,\alpha_{i})$, where $(x,t)\in\mathrm{conf}_{n}$, $\gamma$ is a homotopy class of paths from this point to $\mathrm{ord}$, and $\alpha_{i}$, for $1\leq i\leq n$, is an $n$ tuple of points in $X$, where $n=|x|$. Here the braid group action is such that $j$th standard generator of $B_{n}$, $\sigma_{j}$, sends $(x,t,\gamma,\alpha_{i})$ to $(x,t,\sigma_{j}\gamma,\sigma_{j}(\alpha_{i}))$ using § A.2 and § A.2 to get a left action of $B_{n}$ on such $\gamma$, and using the left action of the braid group on $X^{n}$ from § A.1.
@@
The key construction we use for Hurwitz spaces for braided sets is the following cutting construction. The idea for this construction is that if we start with a configuration in $0,t^{\prime}]\times[0,1]$ which never contains points in the vertical line $t\times[0,1]$, we can cut it into two configurations by cutting along $t\times[0,1]$.
@@
Let $t\in[0,\infty)$, and consider the open set $U(t)$ of $\operatorname{hurbig}^{X}$ consisting of $(x,t^{\prime},\gamma,\alpha_{i})$ such that $t^{\prime}>t$ and no points of the configuration $x$ lie on $t\times[0,1]$. Then there is a continuous map $U(t)\to\operatorname{hurbig}^{X}\times\operatorname{hurbig}^{X}$ which we call ‘cutting at $t$’, given as follows: We first cut the configuration on $0,t^{\prime}]$ into two configurations $x^{\prime},x^{\prime\prime}$ on $[0,t]$ and $0,t^{\prime}-t]\cong[t,t^{\prime}]$. Choose a homotopy class of paths $\phi$ from $x$ to $\mathrm{ord}$ with a representative that doesn’t cross $t\times[0,1]$. Using the braid group action, the tuple $(x,t^{\prime},\gamma,\alpha_{i})$ is equivalent to $(x,t^{\prime},\phi,\beta_{i})$ for some $\beta_{i}$ obtained by acting on the tuple $(x_{i})$ via the element of the braid group given by the class of $\phi\gamma^{-1}$. We can now similarly cut $\phi$ into homotopy classes of paths $\phi^{\prime},\phi^{\prime\prime}$ from the configurations $x^{\prime},x^{\prime\prime}$ to the respective base points, and cut the sequence $\beta_{i}$ into $\alpha_{j}^{\prime},\alpha_{k}^{\prime\prime}$ by unconcatenating it into two sequences corresponding to the number of points in the configuration in $0,t]\times[0,1]$ and $t,t^{\prime}]\times[0,1]$ respectively.
Then the cutting map sends $(x,t^{\prime},\gamma,\alpha_{i})$ to the pair of tuples $(x^{\prime},t,\phi^{\prime},\alpha^{\prime}_{j})$ and $(x^{\prime\prime},t^{\prime}-t,\phi^{\prime\prime},\alpha^{\prime\prime}_{k})$. Since the concatenation map $X^{a}\times X^{b}\to X^{a+b}$ is $B_{a}\times B_{b}$-equivariant, this doesn’t depend on the choice of $\phi$.